Krystal plans to save $500 at the end of Year 1, $600 at the end of Year 2, and $800 at the end of Year 3. If she earns 2.8 percent on her savings, how much money will she have saved at the end of Year 3?
Answer #1

Future value of year 1 cash flow = 500 (1 + 0.028)2 = 528.392

Future value of year 2 cash flow = 600 (1 + 0.028)1 = 616.8

Future value of year 3 cash flow = 800 (1 + 0.028)0 = 800

Future value = 528.392 + 616.8 + 800

Future value = $1,945.19

She will have saved $1,945.19